# ProSparse: Introducing and Enhancing Intrinsic Activation Sparsity within Large Language Models

### v2

@@ 5 Conclusion
In this work, we propose ProSparse, an effective ReLUfication method for introducing and enhancing intrinsic activation sparsity from non-ReLU LLM checkpoints without impairing performance. Extensive experiments not only demonstrate the advantage of ProSparse over existing methods but also reveal its practical values in inference acceleration with both approximate algorithms and accurate algorithms. Deeper analyses of the sparsity trends, dataset-wise sparsity, and layer-wise sparsity indicate the reasonableness of progressive sparsity regularization in smoothing training dynamics, the ProSparse’s high-sparsity bias for more practical instruction tuning datasets, and the effectiveness of activation threshold shifting in alleviating layer-wise sparsity imbalance.
Limitations
Firstly, the scales of LLMs involved in this work, including 7B and 13B, are relatively limited. For more comprehensive studies, both small-scale models (e.g., 2B or less) and huge-scale models (e.g., 70B or more) are supposed to be tested.
Next, we only focus on the sparsity-based acceleration of step (2) and step (3) of FFN, leaving a considerable ratio of LLM computation unoptimized. Actually, there already exist preliminary works in the sparsification of the attention layers Shen et al. (2023); Wortsman et al. (2023). For future works, we will continue to explore how to introduce and enhance sparsity in FFN step (1) as well as the attention layer.
Finally, although the step number of training stages can be easily determined through the validation loss or performance, it is still quite empirical and expensive to search for the best group of peak regularization factors. In the future, we will try to find the quantitative relationships between the sparsity and the $L_{1}$ regularization factor. In this way, it will be much easier to locate proper factors according to the target sparsity.
Ethics Statement
The authors of this work declare that they have no conflict of interest. Besides, no animal or human being is involved as the study objective in any part of this article.
Moreover, we use the open-source LLM LLaMA2 Touvron et al. (2023b) in this paper, which is licensed under the Meta LLAMA 2 Community License. Our usage purpose is only limited to academic research and does not violate the Acceptable Use Policy for the Llama Materials [^7] https://llama.meta.com/use-policy.
@@ Appendix A Extended Related Works of $L_{1}$ Regularization
$L_{1}$ regularization is a classical technique widely used in statistical learning such as linear regression Tibshirani (1996); Hastie et al. (2009). With the advent of deep learning, researchers also explore paradigms of applying $L_{1}$ regularization to neural networks. One prominent usage is model pruning Cheng et al. (2017). Specifically, a term of loss calculated as the $L_{1}$ norm of the sparsification target is added to the optimization target function to prompt sparse weights for faster computation. This has helped acceleration in various conventional neural networks Han et al. (2015b); Zhao et al. (2016); Wen et al. (2016); Scardapane et al. (2017); Ma et al. (2019); Wang et al. (2019) as well as Transformer-based models Zhu et al. (2021); Prasetyo et al. (2023). Inspired by these works, some researchers also try to adopt $L_{1}$ regularization for activation sparsity, mainly in ReLU-activated convolutional networks Georgiadis (2019); Kurtz et al. (2020) and Transformer Li et al. (2022).
To the best of our knowledge, ProSparse is the first work using a dynamic $L_{1}$ regularization factor for prompting activation sparsity in neural networks. By contrast, a majority of the former works adopt fixed factors. For more adaptive control, some of them introduce group regularization Yuan and Lin (2006), namely using different factors for different parameter groups. Nevertheless, without dynamic factors, these paradigms can cause a substantial shift in activation distribution and thus potentially risk performance degradation. The work most related to ProSparse is Wang et al. (2019), which introduces incremental regularization factors that change for different parameter groups at each iteration. While they focus on the pruning of convolutional networks, ProSparse handles a distinct scenario of prompting activation sparsity in Transformer-based LLMs and adopts a completely different strategy consisting of a progressively incremental factor.
@@ Appendix B Extented Introduction of Approximate Acceleration Algorithms
Existing approximate algorithms are mostly dependent on activation predictors, which are small neural networks to predict the intermediate activations $\mathbf{x}_{1}$ based on the input hidden states $\mathbf{x}$ Liu et al. (2023); Song et al. (2023). If one element at a specific position of $\mathbf{x}_{1}$ is predicted to be zero, then all the computations associated with this position can be saved with little or no hardware resources allocated. This is the key mechanism with which approximate algorithms can often reach a high hardware utilization rate and speedup ratio.
Nevertheless, such a predictor-dependent acceleration effect is largely dependent on the performance of the pre-trained activation predictors. For example, a typical bad case is that an actually activated element in $\mathbf{x}_{1}$ is predicted to be inactivated. This can bring about negative results including unwise hardware resource allocation and erroneously ignored intermediate logits, which limits the practical speedup ratios and even causes inference inaccuracies. Therefore, a sparse LLM can gain more benefits from approximate algorithms if its activation distribution is more predictable by the activation predictor.
@@ Appendix D Training Details of Activation Predictors
Following Deja Vu Liu et al. (2023), the predictor is a two-layer FFN, composed of two linear projection layers with a ReLU activation in between them. Notably, as each layer of a sparse LLM has different activation distributions, we should introduce the same number of predictors as that of Transformer layers. For predictor training, we first collect training data with about 400,000 pairs of input hidden states $\mathbf{x}$ and intermediate activations $\mathbf{x}_{1}$ at the corresponding layer. Next, we train the predictor on 95% pairs with the binary cross entropy loss and compute the predictability metrics on the remaining 5% pairs. We reserve the checkpoint with the highest recall to ensure the best inference accuracy with the least falsely ignored activations.
@@ Appendix E Performance on Independent Benchmarks
Appendix F Important Hyperparameters
We provide the important hyperparameters for ProSparse training, as shown in Table 4. Note that the peak regularization factors of two contiguous stages can be set to the same value to introduce an extra constant-factor stage, mainly for stability requirements. All the baseline models are trained with the same number of tokens and the same mixed pre-training dataset as ProSparse. We use a cosine annealing learning rate scheduler throughout the training process and the peak learning rates are $53e-$ 5 and $55e-$ 5 for ProSparse-7B and ProSparse-13B respectively. The context length is 4,096 for all settings.
All the 7B models are trained on 8 A100 GPUs for about 10 days. All the 13B models are trained on 32 A100 GPUs for about 20-30 days. The LLMs of each method involved in this paper are trained once due to the formidable training costs.
Appendix G Effect of Different Thresholds in Activation Threshold Shifting
As mentioned in Section 3.2, the threshold $T$ is an important hyper-parameter in activation threshold shifting, the last step of ProSparse. In the overall experimental results, we choose $T=0.01$ for both ProSparse-7B and ProSparse-13B to balance the sparsity and performance. For comprehensiveness, we list the results under other thresholds in Table 6. As can be observed, a small $T$ results in a quite limited sparsity improvement compared with the version without activation threshold shifting, while a large $T$ can cause non-negligible performance degradation. Therefore, we choose $T=0.01$ to strike a balance.
Appendix H Effect of Different Biases in Shifted ReLU
In the overall experimental results, we set $b=0.1$ for Shifted ReLU-7B, which guarantees the best average performance. In this section, we list the results of Shifted ReLU-7B under different bias $b$ in Table 7. With the increase of bias $b$, the final sparsity of Shifted ReLU models increases marginally, but the performance also suffers from consistent drops. Therefore, the choice of $b$ in Shifted ReLU is also a trade-off between sparsity and performance.
Appendix I Evaluation Details
For PIQA, SIQA, HellaSwag, WinoGrande, COPA, BoolQ, LAMBADA, TyDi QA, and AGI-Eval, we obtain the predicted answers based on maximized perplexity. For GSM8K, MMLU, and BBH, the predicted answers are directly generated.
@@ Appendix I Evaluation Details
Stage Number $i$ $\lambda_{i}$ $T_{i}$ Accumulated Token (B) Stage Number $i$ $\lambda_{i}$ $T_{i}$ Accumulated Token (B)
@@
4 $15e-$ 1 16,000 33.55 4 $22e-$ 2 15,000 125.83
5 $15e-$ 1 16,500 34.60 5 $22e-$ 2 16,000 134.22
Table 4: The important hyperparameters for training ProSparse models. For simplicity, the 0th stage refers to the continual training in activation function substitution. The 1st stage is the warmup stage with a fixed regularization factor $\lambda$. The remaining stages are incremental stages with an increasing $\lambda$.
Vanilla ReLU-7B 18.29 24.33 78.35 50.36 74.29 64.64 86.00 79.42 69.55 70.68
Shifted ReLU-7B 17.07 23.92 78.40 50.31 73.80 63.93 84.00 78.84 69.09 71.59
Fixed $L_{1}$-7B 17.07 20.64 76.06 44.32 68.30 63.38 78.00 52.08 65.01 49.09
$*{}^{*}$ 16.46 22.48 75.79 43.50 71.08 64.09 77.00 62.48 67.73 59.77
$*{}^{*}$ 25.61 32.44 77.04 45.14 75.91 68.67 82.00 79.27 71.08 52.27
Table 5: The performance (%) on each independent benchmark.
@@
Sparsity Generation Reasoning Comprehension
$*{}^{*}$ 88.11 19.47 66.29 63.33 12.74 45.21 33.59 27.55 38.31
@@
Table 6: The sparsity (%) and performance (%) under different thresholds $T$ of activation threshold shifting.
Sparsity Generation Reasoning Comprehension
Shifted ReLU-7B $b=0.1$ 69.59 20.50 70.09 73.17 13.87 48.54 35.20 27.94 41.33
Shifted ReLU-7B $b=0.3$ 70.34 21.05 70.70 72.03 14.33 48.03 34.11 27.58 41.12
Shifted ReLU-7B $b=0.5$ 71.64 20.59 70.25 73.69 13.27 46.55 35.29 27.21 40.98
Shifted ReLU-7B $b=1.0$ 74.81 18.05 70.96 70.64 11.83 46.07 35.48 27.46 40.07
Table 7: The sparsity (%) and performance (%) under different biases $b$ of Shifted ReLU.

### v3

@@ 5 Conclusion
In this work, we propose ProSparse, an effective ReLUfication method for introducing and enhancing intrinsic activation sparsity from non-ReLU LLM checkpoints with comparable performance. Extensive experiments not only demonstrate the effectiveness of ProSparse but also reveal its practical values in inference acceleration with both approximate algorithms and accurate algorithms. Deeper analyses concerned with the control of activation sparsity, the rationality of progressive sparsity regularization, the training dynamics with regularization, SFT on ProSparse models, and the sparsity distributions are performed to substantiate the practicality of ProSparse and provide valuable insights.
@@ Appendix A Extended Related Works of $L_{1}$ Regularization
$L_{1}$ regularization is a classical technique widely used in statistical learning such as linear regression [70, 27]. With the advent of deep learning, researchers also explore paradigms of applying $L_{1}$ regularization to neural networks. One prominent usage is model pruning [11]. Specifically, a term of loss calculated as the $L_{1}$ norm of the sparsification target is added to the optimization target function to prompt sparse weights for faster computation. This has helped acceleration in various conventional neural networks [26, 89, 78, 63, 47, 74] as well as Transformer-based models [93, 57]. Inspired by these works, some researchers also try to adopt $L_{1}$ regularization for activation sparsity, mainly in ReLU-activated convolutional networks [23, 38] and Transformer [43].
To the best of our knowledge, ProSparse is the first work using a dynamic $L_{1}$ regularization factor for prompting activation sparsity in neural networks. By contrast, a majority of the former works adopt fixed factors. For more adaptive control, some of them introduce group regularization [84], namely using different factors for different parameter groups. Nevertheless, without dynamic factors, these paradigms can cause a substantial shift in activation distribution and thus potentially risk performance degradation. The work most related to ProSparse is IncReg [74], which introduces incremental regularization factors that change for different parameter groups at each iteration. While they focus on the pruning of convolutional networks, ProSparse handles a distinct scenario of prompting activation sparsity in Transformer-based LLMs and adopts a completely different strategy consisting of a progressively incremental factor.
@@ Appendix B Extented Introduction of Approximate Acceleration Algorithms
Existing approximate algorithms are mostly dependent on activation predictors, which are small neural networks to predict the intermediate activations $\mathbf{x}_{1}$ based on the input hidden states $\mathbf{x}$ [44, 66]. If one element at a specific position of $\mathbf{x}_{1}$ is predicted to be zero, then all the computations associated with this position can be saved with little or no hardware resources allocated. This is the key mechanism with which approximate algorithms realize acceleration.
Nevertheless, such a predictor-dependent acceleration effect is largely dependent on the performance of the pre-trained activation predictors. For example, a typical bad case is that an actually activated element in $\mathbf{x}_{1}$ is predicted to be inactivated. This can bring about negative results including unwise hardware resource allocation and erroneously ignored intermediate logits, which limits the practical speedup ratios and even causes inference inaccuracies. Therefore, a sparse LLM can gain more benefits from approximate algorithms if its activation distribution is more predictable by the predictors.
@@ Appendix D Training Details of Activation Predictors
Following Deja Vu [44], the predictor is a two-layer FFN, composed of two linear projection layers with a ReLU activation in between them. Notably, as each layer of a sparse LLM has different activation distributions, we should introduce the same number of predictors as that of Transformer layers. For predictor training, we first collect training data with about 400,000 pairs of input hidden states $\mathbf{x}$ and intermediate activations $\mathbf{x}_{1}$ at the corresponding layer. Next, we train the predictor on 95% pairs with the binary cross entropy loss and compute the predictability metrics on the remaining 5% pairs. We reserve the checkpoint with the highest recall to ensure the best inference accuracy with the least falsely ignored activations.
@@ Appendix E Performance on Independent Benchmarks
Table 4: The performance (%) on each independent benchmark.
ProSparse-7B 16.46 22.48 75.79 43.50 71.08 64.09 77.00 62.48 67.73 59.77
ProSparse-13B 25.61 32.44 77.04 45.14 75.91 68.67 82.00 79.27 71.08 52.27
MiniCPM-1B 41.46 32.24 74.10 48.87 67.27 59.12 69.00 69.30 58.18 55.23
ProSparse-1B 48.78 33.98 73.99 47.54 68.46 59.75 73.00 74.92 53.74 53.41
ProSparse-1B 50.61 33.47 74.16 47.54 68.40 59.75 72.00 75.08 53.74 53.36
Appendix F Ablation Studies of Progressive Sparsity Regularization
Here we provide more detailed experimental results about the ablation of progressive sparsity regularization, as shown in Table 5. Note that for ablation settings, we keep the regularization factor constant without progressive increase. More specifically, the whole training process consists of three steps: the activation function substitution, continual training with a constant regularization factor, and activation threshold shifting.
Table 5: Ablation study results about progressive sparsity regularization, the second step of ProSparse. $\lambda$ refers to the constant regularization factor in the second stage of ablation settings.
Scale $\lambda$ Average Sparsity (%) Average Performance (%) Scale $\lambda$ Average Sparsity (%) Average Performance (%)
7B $25e-$ 2 85.95 37.83 13B $21e-$ 2 86.23 43.87
7B $11e-$ 1 88.62 36.34 13B $22e-$ 2 88.96 42.85
7B $15e-$ 1 93.15 33.86 13B $25e-$ 2 92.65 40.13
7B ProSparse 89.32 38.46 13B ProSparse 88.80 44.90
Appendix G Important Hyperparameters
We provide the important hyperparameters for ProSparse training, as shown in Table 6 and Table 7. Note that the peak regularization factors of two contiguous stages can be set to the same value to introduce an extra constant-factor stage, mainly for stability requirements. For ProSparse-7B and ProSparse-13B, We use a cosine annealing learning rate scheduler throughout the training process and the peak learning rates are $53e-$ 5 and $55e-$ 5 for 7B and 13B respectively. For ProSparse-1B, we use exactly the same hyper-parameter settings as MiniCPM-1B [34] except for the $L_{1}$ regularization. After pre-training on the language modeling dataset with the paradigm of ProSparse, following the original practice, we add an extra decay stage (mixed data with a decreasing learning rate) and an SFT stage (only instruction tuning data with a fixed learning rate). Each of the additional stages has a constant regularization factor. The context length is 4,096 for all settings.
All the 7B models are trained with the AdamW optimizer on 8 A100 80GB GPUs for about 10 days. All the 13B models are trained on 32 A100 80GB GPUs for about 20-30 days. The LLMs of each method involved in this paper are trained once due to the formidable training costs.
Table 6: The important hyperparameters for training ProSparse-7B and ProSparse-13B. For simplicity, the 0th stage refers to the continual training in activation function substitution. The 1st stage is the warmup stage with a fixed regularization factor $\lambda_{1}$. The remaining stages are incremental stages with an increasing factor.
@@ Appendix G Important Hyperparameters
Stage Number $i$ $\lambda_{i}$ $T_{i}$ Accumulated Tokens (B) Stage Number $i$ $\lambda_{i}$ $T_{i}$ Accumulated Tokens (B)
@@
4 $12e-$ 1 16,000 33.55 4 $22e-$ 2 15,000 125.83
5 $12e-$ 1 16,500 34.60 5 $22e-$ 2 16,000 134.22
Table 7: The important hyperparameters for training ProSparse-1B. Compared with the other two settings, we follow the original practice of training MiniCPM-1B [34], appending an extra decay stage and an SFT stage. Note that each of the additional stages has a constant regularization factor.
ProSparse-1B
Stage Number $i$ $\lambda_{i}$ $T_{i}$ Accumulated Tokens (B)
0 0 10,000 49.15
1 $31e-$ 3 15,000 73.73
2 $35e-$ 3 20,000 98.30
3 $35e-$ 3 25,000 122.88
4 $25e-$ 2 35,000 172.03
decay $25e-$ 2 (fixed) 95,000 466.94
SFT $21e-$ 2 (fixed) 101,000 473.02
Appendix H Dataset-Wise Sparsity Distribution
Table 8: The average sparsity (%) on our mixed training dataset (denoted as “Mixed”) and its components, divided into language modeling datasets and instruction tuning datasets.
P3 Instructions Instructions
ProSparse-7B 88.11 88.20 83.30 84.24 91.23 97.94 96.74 90.76 93.00 95.71
ProSparse-13B 87.97 87.50 81.64 83.06 92.45 98.41 97.54 91.65 92.92 96.40
ProSparse-1B 86.25 86.84 82.72 83.23 88.50 89.83 83.36 83.93 90.16 90.70
ProSparse-1B 87.89 88.44 84.71 85.17 89.93 91.01 85.36 85.82 91.36 91.76
Despite the satisfactory average sparsity, there still exist gaps between the mixed training dataset and the actual input texts that the model will encounter in real-life applications. To investigate the sparsity of our model under different scenarios, we compute the sparsity on each component of the mixed dataset respectively.
As demonstrated in Table 8, the sparse LLMs obtained through ProSparse have a pronounced property of inconsistent dataset-wise sparsity. Concretely, the sparsity on instruction tuning datasets is significantly higher than those on language modeling datasets (i.e., StarCoder, Wikipedia, and Pile). Considering the contents of datasets, we come to the following assumption: the more formatted a dataset is, the less hybrid knowledge is needed for generation, and thus the $L_{1}$-regularized models can achieve higher activation sparsity with fewer neurons activated. Plain text datasets including Wikipedia and Pile have the lowest sparsity, followed by the more formatted code dataset StarCoder. Among instruction tuning datasets, QA datasets (e.g., Choice P3) with the most monotonic input-output formats obtain the highest sparsity. By contrast, the sparsity is relatively lower on UltraChat and Flan, covering general dialogues and a wide variety of tasks respectively. Notably, dialogues and tasks with formatted instructions cover a majority of input contents of conversational AI, the mainstream application form of LLMs. Such higher sparsity on instruction tuning data will endow ProSparse with more significant practical values.
Appendix I Layer-Wise Sparsity Distribution
Figure 4: The layer-wise sparsity of ProSparse models. The marker “” denotes the settings without activation threshold shifting.
Appendix J Effect of Different Thresholds in Activation Threshold Shifting
Table 9: The sparsity (%) and performance (%) under different thresholds $t$ of the activation threshold shifting step.
@@ Appendix I Layer-Wise Sparsity Distribution
Sparsity Generation Reasoning Comprehension Performance
ProSparse-7B 88.11 19.47 66.29 63.33 12.74 45.21 33.59 27.55 38.31
@@
As mentioned in Section 3.2, the threshold $t$ is an important hyper-parameter in activation threshold shifting, the last step of ProSparse. In the overall experimental results, we choose $t=0.01$ for both ProSparse-7B and ProSparse-13B to balance the sparsity and performance. Here we list the results under other thresholds in Table 9. As can be observed, a small $t$ results in a quite limited sparsity improvement compared with the version without activation threshold shifting, while a larger $t$ can cause more performance degradation. Therefore, we choose $t=0.01$ to strike a balance.
Appendix K Evaluation Details
For evaluation benchmarks including PIQA, SIQA, HellaSwag, WinoGrande, COPA, BoolQ, LAMBADA, TyDi QA, and AGI-Eval, we obtain the predicted answers based on maximized perplexity. Specifically, the predicted answer to a given question corresponds to the candidate that produces the lowest perplexity when it is concatenated to the question. For GSM8K, MMLU, and BBH, the predicted answers are determined by the option numbers directly generated by the models.